# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v1

@@ 5 Conclusion
We propose DyGMamba, an efficient CTDG representation learning model that can capture long-term temporal dependencies. DyGMamba first leverages a node-level SSM to encode long sequences of historical node interactions. It then employs a time-level SSM to learn edge-specific temporal patterns. The learned patterns are used to select the critical part of the encoded temporal information. DyGMamba achieves superior performance on dynamic link prediction, and moreover, it shows high efficiency and strong scalability compared with previous CTDG methods, implying a great potential in modeling huge amounts of temporal information with a limited computational budget.
Acknowledgement
The authors would like to thank the Federal Ministry of Education and Research and the state governments (www.nhr-verein.de/unsere-partner) for supporting this work/project as part of the joint funding of National High Performance Computing (NHR). Part of the experiments of this work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. This work has also been partially funded by the Munich Center for Machine Learning and supported by the Federal Ministry of Education and Research and the State of Bavaria. Zifeng Ding acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under ELISE Grant Agreement No 951847.
Appendix A Complexity Analysis between DyGMamba and DyGFormer
Given a batch size of $B$, a sequence length of $\rho/p$, a patch feature dimension size $4d$ and an SSM dimension $d_{SSM}$, we compute DyGMamba’s required computation for sequence encoding as follows. DyGMamba consists of a node-level SSM block and a time-level SSM block. For the node-level SSM block, the complexity is $O(B\cdot(\rho/p)\cdot 4d\cdot d_{SSM})$. For the time-level SSM block, the complexity is $O(B\cdot k\cdot\gamma d\cdot d_{SSM})$. Note that DyGMamba processes source and destination nodes’ sampled neighbors separately in the node-level SSM block. By considering all of these conditions, we can write DyGMamba’s complexity as $O(Bd_{SSM}d(8(\rho/p)+k\gamma))$. In practice, we choose $d_{SSM}$ as the default value of $16$. We also choose $\gamma$ as $0.5$ for all experiments. Based on this, we can further write DyGMamba’s complexity as $O(128Bd(\rho/p)+8Bdk)$. Recall that $k$ in practice is small, e.g., 10, we can thus write DyGMamba’s complexity as $O(\rho/p)$. DyGFormer concatenate source and destination nodes’ sampled neighbors into one sequence, so its complexity for sequence encoding is $O(B\cdot 4d\cdot(2\rho/p)^{2}+B\cdot(4d)^{2}\cdot(2\rho/p))=O(16Bd((\rho/p)^{2}+2\rho/p))=O((\rho/p)^{2})$. To this end, we can conclude that DyGMamba is much more scalable than DyGFormer when they are used to model long node interaction sequences.
Appendix B CTDG Dataset Details
We present the dataset statistics of all considered CTDG datasets in Table 5. All the datasets in our experiments are taken from (Yu et al., 2023). Please refer to it for detailed dataset descriptions.
@@ Appendix B CTDG Dataset Details
Appendix C Baseline Details
@@
C.1 Baselines Not Designed for Long-Range Temporal Information Propagation
@@
  • DyRep (Trivedi et al., 2019): DyRep updates node representations as events appear. It designs a two-time scale deep temporal point process approach for source and destination nodes and couple the structural and temporal components with a temporal-attentive aggregation module.
@@
Note that TGN uses a memory network to store the whole graph history, making it able to preserve long-range temporal information. However, as disscussed in (Yu et al., 2023), it faces a problem of vanishing/exploding gradients, preventing it from optimally capturing long-term temporal dependencies. EdgeBank can also preserve a very long graph history, but we can observe from the experimental results (Table 1, 8) that without learnable parameters, it is not strong enough on long-range temporal dependent datasets.
C.2 Baselines Designed for Long-Range Temporal Information Propagation
@@
  • DyGFormer (Yu et al., 2023): DyGFormer is a Transformer-based CTDG model. It takes the long-term one-hop temporal interactions of source and destination nodes and use a Transformer to encode them. A patching technique is developed to cut the computational consumption and a node co-occurrence encoding scheme is used to exploit the correlations of nodes in each interaction. DyGFormer achieves long-range temporal information propagation by increasing the number of sampled one-hop historical interactions. The patching technique ensures that even with a huge number of sampled interactions, the length of the sequence input into Transformer will not be too long, making it possible to implement DyGFormer with a limited computational budget.
@@
  • CTAN (Gravina et al., 2024): CTAN is deep graph network for learning CTDGs based on non-dissipative ordinary differential equations. CTAN’s formulation allows for a scalable long-range temporal information propagation in C-TDGs because its non-dissipative layer can retain the information from a specific event indefinitely, ensuring that the historical context of a node is preserved despite the occurrence of additional events involving this node.
Appendix D Implementation Details
@@
We train every CTDG model except for CTAN for a maximum number of $200$ epochs. Maximum epochs for CTAN training is $1000$. We evaluate each model on the validation set at the end of every training epoch and adopt an early stopping strategy with a patience of 20. We take the model that achieves the best validation result for testing. We use the implementations [^4] https://github.com/yule-BUAA/DyGLib provided by (Yu et al., 2023) for all baseline models except CTAN. For CTAN, we use its official implementation [^5] https://github.com/gravins/non-dissipative-propagation-CTDGs. All models are trained with a batch size of 200 for fair efficiency analysis. All experiments are implemented with PyTorch (Paszke et al., 2019) on a server equipped with an AMD EPYC 7513 32-Core Processor and a single NVIDIA A40 with 45GB memory. We run each experiment for five times with five random seeds and report the mean results together with error bars.
D.1 Hyperparameter Configurations
@@
For all the baselines except CTAN, please refer to (Yu et al., 2023) for the hyperparameter configurations. For CTAN, we present its hyperparameter configurations in Table 6. We keep its hyperparameters unchanged for all datasets. Note that we set the number of graph convolution layers (GCLs) in CTAN to its maximum, i.e., 5, in order to maximize its performance in capturing long-term temporal dependencies.
@@
Table 6: Hyperparameter configurations of CTAN on all datasets. $\gamma$ here denotes the discretization step size introduced in Gravina et al. (2024), different from the one in DyGMamba.
@@
We report the hyperparameter searching strategy of DyGMamba and the best hyperparameters in Table 7. Note that to achieve fair efficiency comparison with DyGFormer, we fix the length of the input sequence into the node-level SSM to $32$, i.e., $\rho\ \&\ p=32$. The results reported in Table 1, 2, 8, 9 are all achieved by DyGMamba with $\rho\ \&\ p=32$. In practice, we can decrease $p$ to have a better performance given more computational resources (as discussed in Sec. 4.3). DyGMamba keeps the embedding size as same as DyGFormer on all datasets, i.e., $d_{N}=d_{E}=172,d_{T}=100,d_{F}=50$. We also set $\gamma=0.5$ for all experiments of DyGMamba. The dimension of SSMs $d_{SSM}=16$ remains the default value of mamba SSM’s official repository [^6] https://github.com/state-spaces/mamba.
Table 7: DyGMamba hyperparameter searching strategy. The best settings are marked as bold.
Dataset Dropout $\rho\ \&\ p$ $k$
@@
Appendix E Negative Edge Sampling Strategies during Evaluation
We justify why we do not do historical NSS for inductive link prediction. As described in Poursafaei et al. (2022), historical NSS focuses on sampling negative edges from the set of edges that have been observed during previous timestamps but are absent in the current step. In the setting of inductive link prediction, models are asked to predict the links between the nodes unseen in the training dataset. This means when doing historical NSS, models only need to care about the previously observed edges in the test set (or validation set during validation) for choosing negative edges. This makes historical NSS as sames as inductive NSS in the inductive link prediction, where inductive NSS samples negative edges that have been observed only in the test set, but not training set. Empirical results shown in Appendix C.2 Table 13 and 14 of Yu et al. (2023) also prove that, there is no difference between historical and inductive NSS in inductive link prediction. So we omit the results of historical NSS in our paper.
Appendix F AUC-ROC Results
Table 8 and 9 presents the AUC-ROC results of all baselines and DyGMamba. We have similar observations as the AP results shown in Table 1 and 2. DyGMamba still demonstrates superior performance and can achieve the best average rank under any NSS setting in both transductive and inductive link prediction.
@@ Appendix F AUC-ROC Results
Appendix G Dynamic Node Classification
@@
We follow (Rossi et al., 2020; Xu et al., 2020; Yu et al., 2023) to conduct dynamic node classification by estimating the state of a node in a given interaction at a specific timestamp. A classification MLP is employed to map the node representations as well as the learned temporal patterns to the labels. AUC-ROC is used as the evaluation metric and we follow the dataset splits introduced in (Yu et al., 2023) ($70\%15\%/15\%$ for training/validation/testing in chronological order) for node classification. Table 10 shows the node classification results on Wikipedia and Reddit (the only two CTDG datasets for dynamic node classfication), we observe that DyGMamba can achieve the best average rank, showing its strong performance. Note that both Wikipedia and Reddit are not long-range temporal dependent datasets, therefore we do not include this part into the main body of the paper. Nonetheless, DyGMamba’s great results on these datasets further prove its strength in CTDG modeling, regardless of the type of the dataset (whether long-range temporal dependent or not).
@@
Appendix H Efficieny Analysis Complete Results
@@
H.1 Efficiency Statistics for all baselines
Table 11: Efficiency statistics for all baselines. EdgeBank is non-parameterized and not a machine learning model so we omit it here. # Params means number of parameters (MB). Time and Mem denote per epoch training time (min) and GPU memory (GB), respectively.
@@ H.1 Efficiency Statistics for all baselines
H.2 Total Training Time Comparison among DyGFormer, CTAN and DyGMamba
@@
H.3 Modeling an Increasing Number of Temporal Neighbors with Limited Total Training Time
@@
To further show DyGMamba’s superior efficiency against baseline methods, we do the following experiments. We train five best performing models (as shown in Table 1) on Enron with a gradually increasing number of temporal neighbors [^7] For CTAN, by number of temporal neighbors we mean the sampler size in each graph convolutional layer, i.e., the size of the sampled temporal neighborhood for each node at a timestamp. and report their performance. The number of sampled neighbors spans from $8$, $16$, $32$, $64$, $128$ to $256$ (Note that these numbers are different from the best hyperparamters reported in (Yu et al., 2023)). We fix the patch size $p$ of DyGFormer and DyGMamba to $1$ in order to maximize their input sequence lengths. We set a time limit of $120$ minutes for the total training time. We let all the experiments finish the complete training process and note down the ones that exceed the time limit. In this way, we not only care about the per epoch training time, but also pay attention to how long it takes for models to converge. The experimental results are reported in Fig. 5. The points marked with crosses ($\times$) mean that the training process cannot finish within the time limit (although we still plot their corresponding performance). We find that only TGN and DyGMamba can successfully converge within the time limit when the number of considered neighbors increases to $256$. DyGMamba can constantly achieve performance gain from modeling more temporal neighbors while TGN cannot. CAWN is extremely time consuming so it cannot finish training within the time limit even when it is asked to model $16$ temporal neighbors. As for the methods designed for long-range temporal information propagation, DyGFormer and CTAN consume much longer total training time than DyGMamba. They fail to converge within $120$ minutes when the number of considered neighbors reaches $128$ and $256$, respectively. We also observe that CTAN’s performance fluctuates greatly with the increasing temporal neighbors, indicating that it is not stable to model a large number of temporal neighbors. This also implies that increasing the amount of historical information will gradually make CTAN harder to converge, which might cause trouble in modeling long range temporal dependent datasets.

### v2

@@ 5 Conclusion
We propose DyGMamba, an efficient CTDG representation learning model that can capture long-term temporal dependencies. DyGMamba first leverages a node-level SSM to encode long sequences of historical node interactions. It then employs a time-level SSM to learn edge-specific temporal patterns. The learned patterns are used to select the critical part of the encoded temporal information. DyGMamba achieves superior performance on dynamic link prediction, and moreover, it shows high efficiency and strong scalability compared with previous CTDG methods, implying a great potential in modeling huge amounts of temporal information with a limited computational budget. Our code is uploaded as supplementary materials. Please refer to it for reproducibility.
Appendix A CTDG Dataset Details
A.1 Real-World Benchmark Datasets
We present the dataset statistics of all considered CTDG datasets in Table 6. All the datasets in our experiments are taken from Yu et al. (2023). We chronologically split each dataset with the ratio of 70%/15%/15% for training/validation/testing. Please refer to it for detailed dataset descriptions.
@@ A.1 Real-World Benchmark Datasets
A.2 Synthetic Datasets
For all of our three synthetic datasets, node and link features are not involved during dataset construction. The construction details are as follows:
  • S1: For each interaction pair $u$ and $v$, their first interaction is at timestamp 00 and the second interaction is generated randomly. Thus, the first time interval is also determined. Starting from the second interval, they follow an increasing trend with a constant velocity of 0.05, i.e., $(t_{i+2}-t_{i+1})-(t_{i+1}-t_{i})$= 0.05. The number of interactions for each node pair is randomly determined and the interaction numbers of all node pairs sum up to 100000.
  • S2: For each interaction pair $u$ and $v$, their first interaction is at timestamp 0 and the second interaction is generated randomly. However, it should be large enough so that the interval will not drop to zero or negative afterwards. Starting from the second interval, they follow a decreasing trend with a constant velocity of 0.05, i.e., $(t_{i+1}-t_{i})-(t_{i+2}-t_{i+1})=$ 0.05. The number of interactions for each node pair is randomly determined and the interaction numbers of all node pairs sum up to 100000.
  • S3: S3 contains 8 periods. In each period, the interactions of each node pair $u$ and $v$ are generated following the same pattern in S1. The number of interactions for each node pair is randomly determined and the interaction numbers of all node pairs sum up to 12000 in the period.
We present the statistics of all synthetic datasets in Table 7. We chronologically split each dataset with the ratio of 70%/15%/15% for training/validation/testing. To better visualize the temporal patterns in each dataset, we pick one pair of interacting nodes and plot the time intervals between neighboring interactions in each dataset in Figure 5. Note that for the periodic dataset S3 (Figure 5c), each of the train, validation and test sets contains at least one start of a new period. This ensures that models have to capture periodic temporal patterns in order to achieve good performance during evaluation, rather than only learning the increasing time intervals as specified in S1.
Furthermore, we provide the information about the numbers of interactions regarding interacting node pairs in Table 8. We show that each node pair is equipped with a substantial number of interactions, meaning that temporal patterns in our synthetic datasets span across long time periods. This encourages models to consider long-term temporal dependencies for better graph reasoning.
Table 7: Synthetic dataset statistics.
Datasets # Nodes # Edges # Timestamps Time Range
S1 7 100,000 96,869 0 - 163241.65
S2 7 100,000 98,004 0 - 1573561.52
S3 7 95,657 95,370 0 - 1771300.40
Table 8: Interaction information of node pairs in synthetic datasets. Complete Dataset includes the numbers of interactions across the whole datasets, including training, validation and testing.
Datasets Avg. # Interactions Min # Interactions Max # Interactions
Training Set S1 1,428.57 1,205 1,663
S2 1,428.57 1,424 1,433
S3 1,367.91 1,364 1,372
Complete Dataset S1 2,010.20 1,879 2,150
S2 2,010.20 1,921 2,097
S3 1,952.18 1,721 2,193
S3 (each period) 244.02 215 274
(a) Synthetic dataset S1.
(b) Synthetic dataset S2.
(c) Synthetic dataset S3.
Figure 5: Time intervals of a node pair $u$, $v$ in synthetic datasets S1, S2 and S3.
Appendix B Baseline Details
@@
B.1 Baselines Not Designed for Long-Range Temporal Information Propagation
@@
  • DyRep (Trivedi et al., 2019): DyRep updates node representations as events appear. It designs a two-time scale deep temporal point process approach for source and destination nodes and couples the structural and temporal components with a temporal-attentive aggregation module.
@@
Note that TGN uses a memory network to store the whole graph history, making it able to preserve long-range temporal information. However, as discussed in Yu et al. (2023), it faces a problem of vanishing/exploding gradients, preventing it from optimally capturing long-term temporal dependencies. EdgeBank can also preserve a very long graph history, but we can observe from the experimental results (Table 1, 12) that without learnable parameters, it is not strong enough on long-range temporal dependent datasets.
B.2 Baselines Designed for Long-Range Temporal Information Propagation
@@
  • DyGFormer (Yu et al., 2023): DyGFormer is a Transformer-based CTDG model. It takes the long-term one-hop temporal interactions of source and destination nodes and uses a Transformer to encode them. A patching technique is developed to cut the computational consumption and a node co-occurrence encoding scheme is used to exploit the correlations of nodes in each interaction. DyGFormer achieves long-range temporal information propagation by increasing the number of sampled one-hop historical interactions. The patching technique ensures that even with a huge number of sampled interactions, the length of the sequence input into Transformer will not be too long, making it possible to implement DyGFormer with a limited computational budget.
@@
  • CTAN (Gravina et al., 2024): CTAN is deep graph network for learning CTDGs based on non-dissipative ordinary differential equations. CTAN’s formulation allows for a scalable long-range temporal information propagation in CTDGs because its non-dissipative layer can retain the information from a specific event indefinitely, ensuring that the historical context of a node is preserved despite the occurrence of additional events involving this node.
Appendix C Implementation Details
@@
We train every CTDG model except for CTAN for a maximum number of 200 epochs. Maximum epochs for CTAN training is 1000. We evaluate each model on the validation set at the end of every training epoch and adopt an early stopping strategy with a patience of 20. We take the model that achieves the best validation result for testing. We use the implementations [^4] https://github.com/yule-BUAA/DyGLib provided by Yu et al. (2023) for all baseline models except CTAN. For CTAN, we use its official implementation [^5] https://github.com/gravins/non-dissipative-propagation-CTDGs. All models are trained with a batch size of 200 for fair efficiency analysis. All experiments are implemented with PyTorch (Paszke et al., 2019) on a server equipped with an AMD EPYC 7513 32-Core Processor and a single NVIDIA A40 with 45GB memory. We run each experiment for five times with five random seeds and report the mean results together with error bars.
C.1 Hyperparameter Configurations on Real-World Datasets
@@
For all the baselines except CTAN, please refer to Yu et al. (2023) for the hyperparameter configurations on real-world datasets. For CTAN, we present its hyperparameter configurations in Table 9. We keep its hyperparameters unchanged for all real-world datasets. Note that we set the number of graph convolution layers (GCLs) in CTAN to its maximum, i.e., 5, in order to maximize its performance in capturing long-term temporal dependencies.
@@
Table 9: Hyperparameter configurations of CTAN on all real-world datasets. $\gamma$ here denotes the discretization step size introduced in Gravina et al. (2024), different from the one in DyGMamba.
@@
We report the hyperparameter searching strategy of DyGMamba on real-world datasets and the best hyperparameters in Table 10. To achieve fair efficiency comparison with DyGFormer, we fix the length of the input sequence into the node-level SSM to $32$, i.e., $\rho\ \&\ p=$ 32. The results reported in Table 1, 2, 12, 13 are all achieved by DyGMamba with $\rho\ \&\ p=$ 32. In practice, we can decrease $p$ to have a better performance given more computational resources (as discussed in Sec. 4.3). DyGMamba keeps the embedding size as same as DyGFormer on all real-world datasets, i.e., $d_{N}=d_{E}=$ 172, $d_{T}=$ 100, $d_{F}=$ 50. We also set $\gamma=$ 0.5 for all experiments of DyGMamba. The dimension of SSMs $d_{SSM}$ = 16 remains the default value of mamba SSM’s official repository [^6] https://github.com/state-spaces/mamba. For all experiments, we set the numbers of layers in both node-level and time-level SSMs as 2.
Table 10: DyGMamba hyperparameter searching strategy on real-world datasets. The best settings are marked as bold.
Datasets Dropout $\rho\ \&\ p$ $k$
@@
Table 11: DyGFormer and DyGMamba hyperparameter searching strategy on synthetic datasets. The best settings are marked as bold.
Models DyGFormer DyGMamba
Datasets $\rho\ \&\ p$ $\rho\ \&\ p$ $k$
S1 {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {30, 10, 5}
S2 {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {30, 10, 5}
S3 {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1} {30, 10, 5}
C.2 Hyperparameter Configurations on Synthetic Datasets
We use the same settings of CTAN on real-world datasets when we experiment it on synthetic datasets. For DyGFormer and DyGMamba, we fix the length of the input sequence into the Transformer and the node-level SSM to 32, i.e., $\rho/p=$ 32. For DyGFormer, we set the hyperparameters except $\rho$ and $p$ to the same default values as on real-world datasets, and search for the best $\rho\ \&\ p$ within {512 & 16, 256 & 8, 128 & 4, 64 & 2, 32 & 1}. For DyGMamba, we search for the best $\rho\ \&\ p$ within the same search range and further search for the best $k$. All the other hyperparameters are set as same as the setting on LastFM. We report the hyperparameter searching strategy as well as the best settings of DyGFormer and DyGMamba on synthetic datasets in Table 11. For all experiments with DyGMamba, we set the numbers of layers in both node-level and time-level SSMs as 2.
Appendix D Negative Edge Sampling Strategies during Evaluation
We justify why we do not do historical NSS for inductive link prediction. As described in Poursafaei et al. (2022), historical NSS focuses on sampling negative edges from the set of edges that have been observed during previous timestamps but are absent in the current step. In the setting of inductive link prediction, models are asked to predict the links between the nodes unseen in the training dataset. This means when doing historical NSS, models only need to care about the previously observed edges in the test set (or validation set during validation) for choosing negative edges. This makes historical NSS the same as inductive NSS in the inductive link prediction, where inductive NSS samples negative edges that have been observed only in the test set, but not training set. Empirical results shown in Appendix C.2 Table 13 and 14 of Yu et al. (2023) also prove that there is no difference between historical and inductive NSS in inductive link prediction. So we omit the results of historical NSS in our paper.
@@ Appendix D Negative Edge Sampling Strategies during Evaluation
Appendix E AUC-ROC Results on Real-World Datasets
Table 12 and 13 presents the AUC-ROC results of all baselines and DyGMamba on real-world datasets. We have similar observations as the AP results shown in Table 1 and 2. DyGMamba still demonstrates superior performance and can achieve the best average rank under any NSS setting in both transductive and inductive link prediction.
@@ Appendix E AUC-ROC Results on Real-World Datasets
Appendix F Dynamic Node Classification
@@
We follow Rossi et al. (2020); Xu et al. (2020); Yu et al. (2023) to conduct dynamic node classification by estimating the state of a node in a given interaction at a specific timestamp. A classification MLP is employed to map the node representations as well as the learned temporal patterns to the labels. AUC-ROC is used as the evaluation metric and we follow the dataset splits introduced in Yu et al. (2023) (70%15%/15% for training/validation/testing in chronological order) for node classification. Table 14 shows the node classification results on Wikipedia and Reddit (the only two CTDG datasets for dynamic node classfication), we observe that DyGMamba can achieve the best average rank, showing its strong performance. Note that both Wikipedia and Reddit are not long-range temporal dependent datasets, therefore we do not include this part into the main body of the paper. Nonetheless, DyGMamba’s great results on these datasets further prove its strength in CTDG modeling, regardless of the type of the dataset (whether long-range temporal dependent or not).
@@
Appendix G Efficieny Analysis Complete Results
@@
G.1 Efficiency Statistics for all baselines
We provide the efficiency statistics for all baselines in Table 15.
Table 15: Efficiency statistics for all baselines. EdgeBank is non-parameterized and not a machine learning model so we omit it here. # Params means number of parameters (MB). Time and Mem denote per epoch training time (min) and GPU memory (GB), respectively. The numbers in this table are the average results of five runs with different random seeds.
@@ G.1 Efficiency Statistics for all baselines
G.2 Total Training Time Comparison among DyGFormer, CTAN and DyGMamba
G.3 Modeling an Increasing Number of Temporal Neighbors with Limited Total Training Time
@@
To further show DyGMamba’s superior efficiency against baseline methods, we do the following experiments. We train five best performing models (as shown in Table 1) on Enron with a gradually increasing number of temporal neighbors [^7] For CTAN, by number of temporal neighbors we mean the sampler size in each graph convolutional layer, i.e., the size of the sampled temporal neighborhood for each node at a timestamp. and report their performance. The number of sampled neighbors spans from 8, 16, 32, 64, 128 to 256 (Note that these numbers are different from the best hyperparameters reported in Yu et al. (2023)). We fix the patch size $p$ of DyGFormer and DyGMamba to $1$ in order to maximize their input sequence lengths. We set a time limit of 120 minutes for the total training time. We let all the experiments finish the complete training process and note down the ones that exceed the time limit. In this way, we not only care about the per epoch training time, but also pay attention to how long it takes for models to converge. The experimental results are reported in Fig. 6. The points marked with crosses ($\times$) mean that the training process cannot finish within the time limit (although we still plot their corresponding performance). We find that only TGN and DyGMamba can successfully converge within the time limit when the number of considered neighbors increases to 256. DyGMamba can constantly achieve performance gain from modeling more temporal neighbors while TGN cannot. CAWN is extremely time consuming so it cannot finish training within the time limit even when it is asked to model 16 temporal neighbors. As for the methods designed for long-range temporal information propagation, DyGFormer and CTAN consume much longer total training time than DyGMamba. They fail to converge within 120 minutes when the number of considered neighbors reaches 128 and 256, respectively. We also observe that CTAN’s performance fluctuates greatly with the increasing temporal neighbors, indicating that it is not stable to model a large number of temporal neighbors. This also implies that increasing the amount of historical information will gradually make CTAN harder to converge, which might cause trouble in modeling long range temporal dependent datasets.